OpenStudy (anonymous):

the foci of an ellipse are (0,+ or - 4) and the length of the major axis is 12. How long is the minor axis?

OpenStudy (amistre64):

a = 12/2 = 6 b= b c = 4 minor axis = b+b; b = sqrt(6^2 - 4^2)

OpenStudy (anonymous):

how does c=4?

OpenStudy (amistre64):

4 up and 4 down from the origin there tells me that the foci are a distance of....well, 4

OpenStudy (amistre64):

c just happens to be a leg of a rt tri and a is the hyp

OpenStudy (anonymous):

is sqrt(6^2-4^2) a formula?

OpenStudy (amistre64):

since a and b are used in the equation of an ellipse; it doesnt follow the usually lettering of the pythag thrm; but the values to get the sides and such remain the same right?

OpenStudy (anonymous):

yea

OpenStudy (amistre64):

to find the value of 'b'; we have the hyp and a leg; pythag it to find b and then b+b = minor axis

OpenStudy (anonymous):

confuuuuuuused.